# An improved parametrized test of general relativity using the IMRPhenomX waveform family: Including higher harmonics and precession

### v1

@@ IV.2.1 Common parameter and multiplying posteriors
If we assume the deviation parameter value is the same across the events, we can combine the evidence to measure that parameter by multiplying individual posteriors for each event. For $N$ events, lets assume the posterior distribution $p(\delta\hat{p}^{(j)}_{i}\mid d^{(j)})$ obtained from the $j$-th data $d^{(j)}$ for the deviation parameter $\delta\hat{p}_{i}$, where $j=1,...,N$. The joint posterior probability is given as,
@@ IV.2.2 Hierarchical approach
×$\times p\left(\mu_{i},\sigma_{i}\mid\{d^{(j)}\}_{j=1}^{N}\right)\;d\mu_{i}\>d\sigma_{i}$ (32)
@@ IV.3 Results and Discussion
We summarize the combined results in Table 3. The hierarchical analysis yields the tightest bound with the -1PN term, $\delta\hat{\varphi}_{-2}=0.01^{+0.29}_{-0.27}\times 10^{-3}$, within a 90% credible interval. The next tightest bounds are found with intermediate parameters, particularly $\delta\hat{b}_{3}=0.04^{+0.18}_{-0.12}\times 0.05$, within the same interval. The widest bound appears with the 3PN log term, $\delta\hat{\varphi}_{6}^{(\ell)}=-0.21^{+1.44}_{-1.56}$. In Table 3, we also report the results from the common parameter analysis under the “Restricted” column, showing consistently tighter 90% credible intervals for $\delta\hat{p}_{i}$ compared to the hierarchical analysis. This is expected as common parameter analysis is a special case of hierarchical analysis works assuming $\sigma_{i}=0$, which can only explore a subspace of entire population of the non-GR parameters. We note that this can lead to biased conclusion if that assumption is not correct. While comparing these two approaches for combining the events, we find significant changes in 0PN, 3.5PN, and several post-inspiral parameters, as reported in terms of $\mathcal{Q}_{\mathrm{GR}}$ in Table 3.
Two widely used approaches in the context of parametrized tests of GR with GW signals are TIGER and FTI. In Appendix B, we compare TIGER results against FTI results reported by LVK. We highlight the results for two GWTC-3 events, GW200225_060421 and GW191204_171526, where the first one shows excellent agreement, and the other has inconsistency in many of the PN terms. The baseline models of these approaches are not the same, and the FTI model neglects the contribution of higher-order modes and the precession effect, which could lead to disagreement. For GWTC-3 events, we have not noticed any significant inconsistency. The combined results reported in Fig. 2 are also comparable to FTI, except for the -1PN parameter. The improvement in -1PN for TIGER results is due to the two GWTC-2 events, GW190814 and GW190412, where the contribution of higher-order modes leads to a narrower posterior, as shown in Fig. 8.
@@ V Summary and conclusion
This TIGER framework with IMRPhenomX models is integrated into the LIGO Algorithm Library (LAL) [46], the parametrized waveforms can be generated using the LALSimulation library  [^1] Available at https://git.ligo.org/lscsoft/lalsuite. The corresponding Bayesian analysis framework is implemented in Bilby TGR package [62]. While compiling this manuscript, this TIGER framework had already been applied in various studies, including testing GR with the GW230529 event [87, 88] and investigating astrophysical environmental effects in tests of vacuum GR [30].
@@ Appendix A Impact of higher harmonics and precession
Figure 6: Same as Fig. 1, but displaying the posterior distributions for all deviation parameters. These TIGER results for the GW200129-like injections are obtained using different waveform models to highlight the impact of higher-order modes and precession effects in parametrized tests, as described in Sec. III and Appendix  A.
Figure 7: Same as Fig. 5, but showing the posterior distributions for all deviation parameters. These TIGER results for the GW190412 and GW190814 events are obtained using different waveform models to highlight the impact of higher-order modes and precession effects in parametrized tests, as discussed in Sec. IV.4 and Appendix  A.
@@ Appendix B Comparison between TIGER and FTI
We present a comparison between our TIGER results and FTI results reported in Ref. [37]. The data for the FTI analysis has been obtained from the LVK data release[92]. Fig. 8 highlights the comparison for two GWTC-3 events. For the GW200225_060421 event, we found that the TIGER and FTI posteriors are largely consistent across all deviation parameters, except for the 3PN logarithmic term. We also highlight results for the GW191204_171526 event, which exhibited the largest inconsistency among the GWTC-3 events, with discrepancies visible in multiple PN terms. Since the baseline models of these two approaches differ, full agreement between the results is not expected. Additionally, the absence of higher-order modes and precession effects in the FTI model could contribute to the inconsistencies if the observed signal contains these components.
We note that the FTI analyses of the GW190814 event were carried out using the SEOBNRv4HM_ROM model to mitigate biases arising from the omission of higher-order harmonics [36]. A comparison with the TIGER results is presented in the bottom row of Fig. 8. The posterior distributions for both methods are consistent at lower PN orders (from -1PN to 1PN). However, significant differences appear at higher PN orders, starting from 2.5PN. Notably, for the 3PN and 3.5PN terms, the GR value lies outside the 90% credible interval in the TIGER analysis, while the FTI posteriors remain consistent with the GR predictions. For all three events, the differences are mostly noticeable for the higher PN terms. Since the baseline model of the FTI framework does not account for precession effects, their omission may potentially lead to inconsistencies.
Figure 8: Comparison of TIGER and FTI results. The top and middle rows correspond to the GWTC-3 events GW200225_060421 and GW191204_171526, respectively. The bottom row shows the comparison for the GW190814 event. The FTI results are obtained from the LVK data release [92]. These results are further discussed in Appendix B. The small horizontal lines indicate the 90% credible intervals of the posteriors, and the gray dashed line at $\delta\hat{\varphi}_{i}=0$ represents the GR prediction.

### v2

@@ IV.2.1 Common parameter and multiplying posteriors
If we assume the deviation parameter value is the same across the events, we can combine the evidence to measure that parameter by multiplying individual posteriors for each event. For $N$ events, lets assume the posterior distribution $p(\delta\hat{p}^{(j)}_{i}\mid d^{(j)})$ obtained from the $j$-th data $d^{(j)}$ for the deviation parameter $\delta\hat{p}_{i}$, where j=1,,Nj=1,...,N. The joint posterior probability is given as,
@@ IV.2.2 Hierarchical approach
$\times p\left(\mu_{i},\sigma_{i}\mid\{d^{(j)}\}_{j=1}^{N}\right)\;d\mu_{i}\>d\sigma_{i}$ (32)
@@ IV.3 Results and Discussion
We summarize the combined results in Table 3. The hierarchical analysis yields the tightest bound with the -1PN term, $\delta\hat{\varphi}_{-2}=-0.01^{+0.15}_{-0.13}\times 0.002$, within a 90% credible interval. The next tightest bounds are found with intermediate parameters, particularly $\delta\hat{b}_{3}=0.16^{+0.17}_{-0.17}\times 0.05$, within the same interval. The widest bound appears with the 3PN log term, $\delta\hat{\varphi}_{6}^{(\ell)}=-0.33^{+1.02}_{-0.85}$. In Table 3, we also report the results from the common parameter analysis under the “Restricted” column, showing consistently tighter 90% credible intervals for $\delta\hat{p}_{i}$ compared to the hierarchical analysis. This is expected as common parameter analysis is a special case of hierarchical analysis works assuming $\sigma_{i}=0$, which can only explore a subspace of entire population of the non-GR parameters. We note that this can lead to biased conclusion if that assumption is not correct. While comparing these two approaches for combining the events, we find significant changes in 0PN, 3.5PN, and several post-inspiral parameters, as reported in terms of $\mathcal{Q}_{\mathrm{GR}}$ in Table 3.
Two widely used approaches in the context of parametrized tests of GR with GW signals are TIGER and FTI. In Appendix B, we compare TIGER results against FTI results reported by LVK. We highlight the results for two O3b events, GW200225_060421 and GW191204_171526, where the first one shows excellent agreement, and the other has inconsistency in many of the PN terms. The baseline models of these approaches are not the same, and the FTI model neglects the contribution of higher-order modes and the precession effect, which could lead to disagreement. For GWTC-3 events, we have not noticed any significant inconsistency. The combined results reported in Fig. 2 are also comparable to FTI, except for the -1PN parameter.
@@ V Summary and conclusion
This TIGER framework with IMRPhenomX models is integrated into the LIGO Algorithm Library (LAL) [49], the parametrized waveforms can be generated using the LALSimulation library  [^2] Available at https://git.ligo.org/lscsoft/lalsuite. There was issue in the -1PN implementation, . The corresponding Bayesian analysis framework is implemented in Bilby TGR package [63]. While compiling this manuscript, this TIGER framework had already been applied in various studies, including testing GR with the GW230529 event [89, 90] and investigating astrophysical environmental effects in tests of vacuum GR [30].
@@
We note that while the IMRPhenomX waveform family has undergone several recent improvements [91, 92, 93, 94], the TIGER results presented in this paper were obtained using IMRPhenomXPHM-MSA (the IMRPhenomXPHM model with the multi–scale analysis (MSA) treatment of spin precession) [43], consistent with the waveform choices in the GWTC-3 catalog [68]. A recent upgrade, IMRPhenomXPHM-SpinTaylor [93], is now being employed for parameter estimation of LVK events observed during the fourth observing run. This version uses the SpinTaylor framework [95] to numerically evolve the spins, providing a more accurate precession evolution compared to the analytical MSA approximation. The TIGER framework works straightforwardly with IMRPhenomXPHM-SpinTaylor, since this model employ the same underlying co-precessing IMRPhenomXHM implementation as in IMRPhenomXPHM-MSA. Further efforts were made to improve the spin-precession sector of the IMRPhenomX waveform family by calibrating the precession angles to numerical relativity, leading to the waveform models IMRPhenomXO4a [91, 92] and IMRPhenomXPNR [94] that have more recently become available. However, those introduce introduce additional modifications to the co-precessing frame, requiring a dedicated TIGER implementation and further validation.
@@ Appendix A Impact of higher harmonics and precession
Figure 6: Same as Fig. 1, but displaying the posterior distributions for all deviation parameters. For some deviation parameters, the posteriors obtained using IMRPhenomXP are significantly broader than those from the other two models. These cases are shown in two consecutive subplots to highlight the difference and to clearly display the narrower posteriors. These TIGER results for the GW200129-like injections are obtained using different waveform models to highlight the impact of higher-order modes and precession effects in parametrized tests, as described in Sec. III and Appendix  A.
Figure 7: Same as Fig. 5, but showing the posterior distributions for all deviation parameters. These TIGER results for the GW190412 and GW190814 events are obtained using different waveform models to highlight the impact of higher-order modes and precession effects in parametrized tests, as discussed in Sec. IV.4 and Appendix  A. For the inspiral parameters, we also show the results obtained with the IMRPhenomPv3HM model, as reported in the GWTC-2 tests of GR paper by LVK [36].
@@ Appendix B Comparison between TIGER and FTI
We present a comparison between our TIGER results and FTI results reported in Ref. [37]. The data for the FTI analysis has been obtained from the LVK data release [99]. Fig. 8 highlights the comparison for two O3b events. For the GW200225_060421 event, we found that the TIGER and FTI posteriors are largely consistent across all deviation parameters, except for the 3PN logarithmic term. We also highlight results for the GW191204_171526 event, which exhibited the largest inconsistency among the GWTC-3 events, with discrepancies visible in multiple PN terms. Since the baseline models of these two approaches differ, full agreement between the results is not expected. Additionally, the absence of higher-order modes and precession effects in the FTI model could contribute to the inconsistencies if the observed signal contains these components.
We note that the FTI analyses of the GW190814 event were carried out using the SEOBNRv4HM_ROM model [100] to mitigate biases arising from the omission of higher-order harmonics [36]. A comparison with the TIGER results is presented in the bottom row of Fig. 8. The posterior distributions for both methods are consistent at lower PN orders (from -1PN to 1PN). However, significant differences appear at higher PN orders, starting from 2.5PN. Notably, for the 3PN and 3.5PN terms, the GR value lies outside the 90% credible interval in the TIGER analysis, while the FTI posteriors remain consistent with the GR predictions. For all three events, the differences are mostly noticeable for the higher PN terms. Since the baseline model of the FTI framework does not account for precession effects, their omission may potentially lead to inconsistencies.
Figure 8: Comparison of TIGER and FTI results. The top and middle rows correspond to the O3b events GW200225_060421 and GW191204_171526, respectively. The bottom row shows the comparison for the GW190814 event. The FTI results are obtained from the LVK data release [99]. These results are further discussed in Appendix B. The small horizontal lines indicate the 90% credible intervals of the posteriors, and the gray dashed line at $\delta\hat{\varphi}_{i}=0$ represents the GR prediction.